### v1

# Frequentist Validity of Epistemic Uncertainty Estimators

Abstract
Decomposing prediction uncertainty into its aleatoric (irreducible) and epistemic (reducible) components is critical for the development and deployment of machine learning systems. A popular, principled measure for epistemic uncertainty is the mutual information between the response variable and model parameters. However, evaluating this measure requires access to the posterior distribution of the model parameters, which is challenging to compute. In view of this, we introduce a frequentist measure of epistemic uncertainty based on the bootstrap. Our main theoretical contribution is a novel asymptotic expansion that reveals that our proposed (frequentist) measure and the (Bayesian) mutual information are asymptotically equivalent. This provides frequentist interpretations to mutual information and new computational strategies for approximating it. Moreover, we link our proposed approach to the widely-used heuristic approach of deep ensembles, giving added perspective on their practical success.
1 Introduction
While machine learning systems are increasingly accurate in tightly-controlled settings, a key challenge to their deployment in more complex, real-world environments is uncertainty awareness—algorithms should be able to accurately report on their level of confidence or reliability in different conditions. Moreover, there are different underlying reasons for uncertainty. Uncertainty may be due to inherent unpredictability in a population with the current feature set (aleatoric or irreducible uncertainty) or instead due to limited training data (epistemic or reducible uncertainty). Distinguishing these types of uncertainty is important in order to take the correct downstream action. High aleatoric uncertainty indicates that the current prediction task is inherently challenging, and we cannot have high confidence in any one prediction unless we are able to collect richer data (not just more data). On the other hand, high epistemic uncertainty signals that collecting more training data in this region is likely to improve the model, which makes it suitable for guiding model development and active learning (Houlsby et al.,, 2011; Gal et al.,, 2017; Smith et al.,, 2023). We focus on this quantity in this work.
A popular measure of epistemic uncertainty is the mutual information (MI) between model parameters and predictions which quantifies the expected reduction in uncertainty from observing new data (Houlsby et al.,, 2011; Depeweg et al.,, 2018). MI is a fundamentally Bayesian quantity that requires access to the posterior distribution over model parameters, which generally renders its exact calculation intractable. As such, substantial effort has gone into developing approximate Bayesian computational techniques (MacKay,, 1995; Graves,, 2011; Welling and Teh,, 2011; Hernandez-Lobato and Adams,, 2015; Gal and Ghahramani,, 2016), but they often require changes to model architectures or training procedures, which may compromise predictive accuracy.
As an alternative approach, our work proposes an estimator of MI based on the bootstrap. In particular, we provide an asymptotic expansion for MI which links it to the Fisher information and shows how epistemic uncertainty differs from frequentist variance. We use this to develop a bootstrap estimator which targets the MI. The estimator is simple—we form bootstrap samples and use them as if they were samples from the posterior distribution, computing the MI accordingly. This requires minimal engineering effort, and does not require any restrictions on the underlying architecture, making it an attractive complement to existing MI estimators. We conduct experiments to showcase the validity of the bootstrapping approach in a deep learning setting and the success of our measure on an active learning task.
Morever, our proposal allows for a natural decomposition of epistemic uncertainty into components arising due to data sampling, and due to stochasticity in the training procedure. This allows us to reason that deep ensembles capture the portion of epistemic uncertainty due to stochastic optimization, which empirically appears to be the majority component, explaining their success. Finally, we provide a proof-of-concept experiment showing how data attribution approaches can be leveraged to provide estimates of epistemic uncertainty.
1.1 Our Contribution
Our contribution can be summarized as follows.
  • Link between mutual information and frequentist Fisher information: We provide an asymptotic expansion for MI and show how it depends not only on the variance in model predictions as measured by the Fisher information, but also on the degree of randomness in the true data generating process.
  • Bootstrap estimation of mutual information: We propose an estimator of MI based on bootstrap. This allows for study between frequentist notions of uncertainty which are based on variance under repeated draws of data, and Bayesian notions which are based on spread in distribution of the parameters.
  • Connection with deep ensembles: We propose a decomposition of our epistemic uncertainty estimator into components arising from training stochasticity and data sampling, and use this to provide insight into the success of deep ensembles.
1.2 Illustration
Figure 1: Teaser Simulation: (left) Training data, true label probability and posterior sample predictions. (center) Predictions from MLEs over bootstrapped datasets. (right) Epistemic uncertainty calculated through Bayesian inference (MCMC) versus our bootstrap based measure.
To provide the key intuition for our method, we provide the following simulation. Consider a binary logistic regression model, $\mathbb{P}(y=1|x,\theta)=1/(1+e^{-\theta_{1}x-\theta_{0}})$. We sample the features $x$ from a standard normal and also impose standard normal priors on $\theta_{0},\theta_{1}$. Fig. 1(left) visualizes the data and the true model along with 100 posterior samples from the true Bayesian posterior, obtained using a Markov chain Monte Carlo (MCMC) algorithm. Next, we calculate maximum likelihood estimates (MLEs) for the parameters across 100 bootstrapped datasets and show these in Fig. 1(middle)—note the remarkable similarity in the posterior and bootstrap samples. Finally, we use these MCMC and bootstrap samples to form our estimate of the epistemic uncertainty as per Algorithm 1 noting the excellent agreement in Fig. 1(right). In the remainder of this work, we develop this connection.
2 Background
2.1 Mutual Information as Epistemic Uncertainty
A popular uncertainty decomposition (MacKay,, 1992; Houlsby et al.,, 2011; Depeweg et al.,, 2018) leverages information-theoretic quantities to propose measures of aleatoric and epistemic uncertainty. Let the random variables $Y$ and $X$ denote the labels and features respectively in a supervised learning setup. Let $\mathcal{D}_{n}$ be the data comprising of $n$ realizations of $(Y,X)$. Further, let $\theta$ be the random variable denoting the parameters of a prediction model. Information theory states that $\mathrm{H}(Y)$ is the uncertainty in $Y$ and defines mutual information as
$\underbrace{\mathrm{I}(Y;\theta|X,\mathcal{D}_{n})}_{\textrm{Epistemic}}=\underbrace{\mathrm{H}(Y|X,\mathcal{D}_{n})}_{\textrm{Total}}-\underbrace{\mathrm{H}(Y|\theta,X,\mathcal{D}_{n})}_{\textrm{Aleatoric}}.$ (1)
The quantification of epistemic uncertainty as the mutual information (MI) between the (test) data and parameters is conceptually pleasing since Eq. 1 reveals it to be the reduction in label uncertainty from knowing the model parameters. This MI is a fundamentally Bayesian object since it requires $\theta$ to be random. In our paper, we will treat the (Bayesian) MI as the targeted measure of epistemic uncertainty and shall investigate various estimators for it.
2.2 Bootstrap
The original bootstrap procedure (Efron,, 1979) was proposed to estimate the variance of an estimator under different draws of the data. Consider a setting where there are no repeated data points. Bootstrap first samples weights from a symmetric multinomial distribution with all event probabilities $1/n$. These weights are used to reweight the original data to create a “bootstrapped dataset” over which the estimator is recomputed. This procedure is replicated $B$ times to obtain $B$ “bootstrapped estimates”. The key intuition behind the bootstrap—the “plug-in” principle—proposes that the distribution of these bootstrapped estimates approximates the distribution of the original estimator under redraws of the data. As such, we can obtain an estimate of the variance of our estimator as the empirical variance of the bootstrapped estimates.
We can also instead draw weights from a symmetric Dirichlet distribution (Rubin,, 1981). We hypothesize that in a deep learning setting, these weights can be more effective at creating bootstrapped datasets since it allows the neural network to still see all the training instances and hence achieve better predictive performance than multinomial weights which omit training points entirely. For our paper, we thus use Dirichlet weights but emphasize that our theorems hold for multinomial weights too (Appendix A).
2.3 Deep Ensembles
Deep ensembles (Lakshminarayanan et al.,, 2017) are a simple but powerful uncertainty quantification technique for deep neural networks (DNNs). A deep ensemble is obtained by training multiple models with the exact same architecture and training data but with different random seeds controlling stochasticity in the training procedure (e.g. initial weights, data shuffling etc.). The deep ensemble model prediction is simply the average of predictions of member models. They have shown tremendous success at a variety of uncertainty quantification tasks (Ashukha et al.,, 2020; Ovadia et al.,, 2019).
3 Algorithm and Theoretical Results
3.1 Problem Setup
We now specialize the previous notation to our exact setting. Consider a $K$-class supervised classification problem with labels $Y\in\{1,...,K\}$. Denote the training data by $\mathcal{D}_{n}=\{(X_{i},Y_{i}),i=1,...,n\}$. We wish to evaluate the epistemic uncertainty in the prediction for the label $Y_{\mathrm{test}}$ given the feature $X_{\mathrm{test}}$. We denote our probability predictions for $Y_{\mathrm{test}}$ for a model parametrized by $\theta\in\Theta$ by $\hat{p}(X_{\mathrm{test}};\theta)$. Further, denote by $\hat{p}_{k}(X_{\mathrm{test}};\theta)$ the prediction probability assigned to the $k^{\textrm{th}}$ class. The model is specified such that $0<\hat{p}_{k}(X_{\mathrm{test}};\theta)<1$ and $\sum_{k=1}^{K}\hat{p}_{k}(X_{\mathrm{test}};\theta)=1$. Finally, we also assume that the model is well-specified, i.e., there exists a $\theta_{0}\in\Theta$ such that $(X_{i},Y_{i})\stackrel{{\scriptstyle iid}}{{\sim}}P_{X}(x)\hat{p}_{y}(x;\theta_{0})=:P_{\theta_{0}}$
In the Bayesian setup, we have a prior $p(\theta)$ over the parameters and we denote by $p(\theta|\mathcal{D}_{n})$ our posterior. We use the notation $\mathbb{E}_{Z}[.]$ to denote expectations taken with respect to the random variable $Z$. As per Sec. 2.1, the epistemic uncertainty is
$\mathrm{I}\left(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n}\right)=\mathrm{H}\left(Y_{\mathrm{test}}|X_{\mathrm{test}},\mathcal{D}_{n}\right)-\mathrm{H}\left(Y_{\mathrm{test}}|\theta,X_{\mathrm{test}},\mathcal{D}_{n}\right)=\mathrm{H}\left(\mathbb{E}_{\theta|\mathcal{D}_{n}}\left[\hat{p}(X_{\mathrm{test}};\theta)\right]\right)-\mathbb{E}_{\theta|\mathcal{D}_{n}}\left[\mathrm{H}(\hat{p}(X_{\mathrm{test}};\theta))\right].$
3.2 Asymptotics of Mutual Information
We begin by deriving the asymptotic behavior of mutual information, connecting it with the Fisher Information and frequentist sampling variation. This connection will allow us to repurpose frequentist methods for estimating variation to instead target the mutual information.
Theorem 1.
Under the assumptions needed for Bernstein von-Mises to hold and uniform integrability of the predicted probabilities ((A1)-(A7) in Appendix A) we have the asymptotic expansion
$\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})=\frac{1}{2n}\sum_{k=1}^{K}\frac{\sigma_{k}^{2}}{\hat{p}_{k}(X_{\mathrm{test}};\theta_{0})}+o_{p}(n^{-1})\qquad\forall\ X_{\mathrm{test}},$ (2)
where $\sigma_{k}^{2}$ is
$\sigma_{k}^{2}=\left[\frac{\partial\hat{p}(X_{\mathrm{test}};\theta_{0})}{\partial\theta}^{T}\mathcal{I}^{-1}(\theta_{0})\frac{\partial\hat{p}(X_{\mathrm{test}};\theta_{0})}{\partial\theta}\right]_{k,k}$
for $\mathcal{I}^{-1}(\theta_{0})$ being the inverse Fisher information of the model at $\theta_{0}$.
Proof Sketch: Our analysis proceeds by Taylor expanding the entropy function about the mean posterior predicted probability. We then apply the functional delta method to the Bernstein von-Mises theorem to obtain the limiting law of our predictions. Recall that the Bernstein von-Mises theorem holds under correct model specification, sufficient smoothness, identifiability, and feasibility of $\theta_{0}$ under the prior (Theorem 10.1 of van der Vaart, (2000)). Finally, we use uniform integrability to obtain the limiting posterior moments. We defer the full proof to Appendix A.
Firstly, the theorem shows that mutual information decreases linearly in the number of samples. Next, it connects mutual information to the Fisher information and perhaps surprisingly shows how the first order asymptotics are governed by the inverse Fisher information divided by the true conditional probabilities.
Better intuition can be gained into the first order asymptotic term by considering the binary classification case in which the first order term simplifies to
$\frac{\left(\partial\hat{p}_{1}(X_{\mathrm{test}};\theta_{0})/\partial\theta\right)^{T}\left(n^{-1}\mathcal{I}^{-1}(\theta_{0})\right)\left(\partial\hat{p}_{1}(X_{\mathrm{test}};\theta_{0})/\partial\theta\right)}{2\hat{p}_{1}(X_{\mathrm{test}};\theta_{0})(1-\hat{p}_{1}(X_{\mathrm{test}};\theta_{0}))}$
as noted in MacKay, (1992). Recall that $n^{-1}\mathcal{I}^{-1}(\theta_{0})$ is the (asymptotic) variance of the MLE. The numerator in the expression above can then be interpreted as the variance of the model prediction at $X_{\mathrm{test}}$. 1/2p(1p)1/2p(1-p) achieves a minima at $1/2$ and tends to infinity at 0 and $1$. The first order term is then directly proportional to the model variance but also inversely proportional to the the ground truth certainty (i.e. how close $\hat{p}_{1}(X_{\mathrm{test}};\theta_{0})$ is to 0 or 1). Thus, we see that this first order term matches the intuition behind epistemic uncertainty; if our predictions are uncertain but the true outcome is certain then we can expect to reduce our uncertainty and hence have high epistemic uncertainty.
3.3 Our Bootstrap Estimator
Theorem 1 suggests that asymptotically correct estimation of the Fisher information and true model probabilities can thus be used to construct first order accurate estimators of mutual information. In light of this, we propose a bootstrap based estimator which we denote by $\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})$.
Our estimation procedure proceeds by first sampling weights $\xi$ from the symmetric Dirichlet with concentration parameter 1. Denote our MLE of $\theta$ over the bootstrapped dataset as $\hat{\theta}_{b}$ (for example, as obtained by minimizing cross-entropy loss). We denote the distribution of this estimate (induced by the randomness in the weights) given the data $\mathcal{D}_{n}$ as $p(\hat{\theta}_{b}|\mathcal{D}_{n})$. Then,
$\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n}):=\mathrm{H}(\mathbb{E}_{\hat{\theta}_{b}|\mathcal{D}_{n}}[\hat{p}(X_{\mathrm{test}};\hat{\theta}_{b})])-\mathbb{E}_{\hat{\theta}_{b}|\mathcal{D}_{n}}[\mathrm{H}(\hat{p}(X_{\mathrm{test}};\hat{\theta}_{b}))].$
Practical calculation is incredibly easy as illustrated by Algorithm 1.
Algorithm 1 $\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})$ estimation
Input: Number of bootstrap replications $B$, training dataset $\mathcal{D}_{n}$
$\Theta_{b}\leftarrow\{\}$
for $b\leftarrow 1$ to $B$ do
  sample $\xi\sim\mathrm{Dir}(1,...,1)$
  $\hat{\theta}_{b}\leftarrow\operatorname*{arg\,max}_{\theta\in\Theta}\sum_{i=1}^{n}\xi_{i}\log(\hat{p}_{Y_{i}}(X_{i};\theta))$
  $\Theta_{b}\leftarrow\Theta_{b}\cup\{\hat{\theta}_{b}\}$
end for
$\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})=\mathrm{H}\left(\frac{1}{B}\sum_{\hat{\theta}_{b}\in\Theta_{b}}\hat{p}(X_{\mathrm{test}};\hat{\theta}_{b})\right)-\frac{1}{B}\sum_{\hat{\theta}_{b}\in\Theta_{b}}\mathrm{H}\left(\hat{p}(X_{\mathrm{test}};\hat{\theta}_{b})\right)$
Output: $\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})$
Theorem 2.
Under the assumptions needed for Bernstein von-Mises to hold, for asymptotic normality of the MLE to hold and uniform integrability of the predicted probabilities ((A1)-(A12) in Appendix A), we have
$\frac{\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})}{\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})}\stackrel{{\scriptstyle P_{\theta_{0}}}}{{\rightarrow}}1\qquad\forall\ X_{\mathrm{test}}.$
Proof Sketch: We proceed as in Theorem 1 but invoke bootstrap limit laws (Theorem 10.16 of Kosorok, (2007)) instead of Bernstein von-Mises to obtain an identical asymptotic expansion for our estimator. We provide the complete proof in Appendix A.
We note that our theorem also holds if we use the classic multinomial weights for bootstrap instead of Dirichlet ones—see Appendix A.
Theorem 2 proves the asymptotic validity of our estimator and justifies its use. Our procedure is model agnostic and can be used even when Bayesian approximations are ad-hoc or entirely unavailable. It thus provides a theoretically justified path to approximating Bayesian information-theoretic quantities using a well studied frequentist procedure.
3.4 Distinguishing Algorithmic Randomness from Statistical Uncertainty
Deep neural networks are generally trained with variants of stochastic gradient descent and a random initialization. As such, the fitted model is random even with the training data fixed. In this section, we show that our bootstrap estimator provides a way to quantify how much of the total variation in the fitted model is due to randomness in the algorithm versus statistical fluctuations in the training set.
In more detail, the parameter estimate can be viewed as a function of the data and the random seed. Thus, for the remainder of this section, we use the notation $\hat{p}(X_{\mathrm{test}};\hat{\theta}_{b})=\hat{p}(\mathcal{D}_{n}^{b},s)$ where $\mathcal{D}_{n}^{b}$ denotes the bootstrapped dataset and $s$ denotes the seed used for training the DNN. Since $\hat{\theta}_{b}$ is deterministic given $(\mathcal{D}_{n}^{b},s)$, $\hat{p}(\mathcal{D}_{n}^{b},s)$ is a deterministic function. Normally, the selection of $s$ is completely independent of everything else and hence $(s|\mathcal{D}_{n}^{b},\mathcal{D}_{n})\stackrel{{\scriptstyle d}}{{=}}s$.
With this notation in place, we propose the following decomposition of the bootstrap epistemic uncertainty estimator
$\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})$
$=$ $\mathrm{H}(\mathbb{E}_{(\mathcal{D}_{n}^{b},s)|\mathcal{D}_{n}}[\hat{p}(\mathcal{D}_{n}^{b},s)])-\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])]+\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])]-\mathbb{E}_{(\mathcal{D}_{n}^{b},s)|\mathcal{D}_{n}}[\mathrm{H}(\hat{p}(\mathcal{D}_{n}^{b},s))]$
$=$ $\underbrace{\mathrm{H}(\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)]])-\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])]}_{=:\mathrm{I}^{\textrm{resampling}}_{b}}+\underbrace{\mathbb{E}_{\mathcal{D}_{n}^{b}|\mathcal{D}_{n}}\left[\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n}^{b},s)])-\mathbb{E}_{s}[\mathrm{H}(\hat{p}(\mathcal{D}_{n}^{b},s))]\right]}_{=:\mathrm{I}^{\textrm{seeds}}_{b}}.$ (3)
$\mathrm{I}^{\textrm{resampling}}_{b}$ can be viewed as an estimator of the epistemic uncertainty arising due to the randomness in data drawing. On the other hand, $\mathrm{I}^{\textrm{seeds}}_{b}$ captures the amount of the entropy difference explained by the variation in the random seeds. That is, this term represents the amount of variation due to the algorithmic randomness. Thus, our proposed estimator can be decomposed into a ‘data sampling’ component and a ‘training randomness’ component.
It is particularly interesting to connect this decomposition with deep ensembles. Notice that the estimate of the mutual information from deep ensembles is
$\mathrm{I}^{\textrm{deep ensemble}}:=$ $\mathrm{H}(\mathbb{E}_{s}[\hat{p}(\mathcal{D}_{n},s)])-\mathbb{E}_{s}[\mathrm{H}(\hat{p}(\mathcal{D}_{n},s))].$
This is parallel to the term $\mathrm{I}^{\textrm{seeds}}_{b}$ above, but with the full data instead of a bootstrap subsample. Thus, the bootstrap estimator above can be viewed as approximately the MI from random seeds plus an additional term capturing variability in the dataset. In the experiments, we will see that the latter is substantially smaller than the former. In this way, $\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})\approx\mathrm{I}^{\textrm{deep ensemble}}$. This link gives a new frequentist motivation for deep ensembles.
4 Experiments
4.1 Validation of the Plug-In Principle
We begin our empirical investigations with validation of the “plug-in” principle—specifically whether the variance under true redraws of the data matches the variance over the bootstrapped datasets.
Since our proposed estimator relies on $\hat{p}(X_{\mathrm{test}};\theta)$, we choose to directly investigate how much variation occurs in $\hat{p}(X_{\mathrm{test}};\theta)$ when we train on multiple datasets drawn from the same population.
Experimental Setup: We use the “The Street View House Numbers” (SVHN) Dataset (Netzer et al.,, 2011) which consist of images of digits of house-numbers, along with the ResNet-18 (He et al.,, 2015) architecture. We defer hyperparameter details to Appendix B.1. We first select at random 50,000 training images from the SVHN dataset (denote this by $\mathcal{D}_{0}$) and train a deep ensemble of 10 models. Let us denote these models by $f^{(1)}_{0},...,f^{(10)}_{0}$ and the deep ensemble model by $\bar{f}_{0}$ (the predictions of the ensemble model are simply the average of constituent model predictions). Next, we select 10 more training datasets each of 50,000 images from the SVHN dataset (denote these by $\mathcal{D}_{1},...,\mathcal{D}_{10}$) such that all the training datasets are disjoint and represent i.i.d. draws of the training data from the same population. We then train a deep ensemble on each of these datasets too (denote the ensemble models by $\bar{f}_{1},...,\bar{f}_{10}$).
Figure 2: (a) Spread under Redraws of Data: Scatter plot of standard deviation in the probability prediction corresponding to the true class for models trained on redraws of the training dataset versus (left) models trained on bootstrapped versions of one dataset and (right) models within a deep ensemble along with Pearson ($\rho$) and Spearman ($r_{s}$) coefficients. See Sec. 4.1 for details. (b) Active Learning: Test accuracy on MNIST for four data acquisition functions. We plot the mean and shade 2 standard deviations across 10 repetitions.
Investigations: To estimate the spread in predictions from $\bar{f}_{0}$, we calculate the standard deviation in the probability predictions corresponding to the true class obtained from $\bar{f}_{1}$ to $\bar{f}_{10}$ for each test instance in the SVHN dataset.
To compare this to the bootstrap estimate, we create 10 bootstrapped datasets from $\mathcal{D}_{0}$ and train a deep ensemble on each of these (denote these by $\bar{f}_{b1},...,\bar{f}_{b10}$). We similarly calculate the standard deviation in the probability predictions corresponding to the true class obtained from $\bar{f}_{b1}$ to $\bar{f}_{b10}$ for each test instance.
Fig. 2(a)(left) shows a scatter plot of the standard deviation under redraws and under bootstrapping for each test instance. As we can see, the two correlate well with a Pearson coefficient of $\rho=0.84$ and a Spearman rank correlation coefficient of $r_{s}=0.86$. The standard deviations under redraws are slightly higher as may be expected due to more variability in training data.
We also use this experimental setup to begin our study into the performance of deep ensembles from a frequentist lens. Specifically, we investigate the spread within a deep ensemble by calculating the standard deviation in the probability prediction corresponding to the true class obtained from $f^{(1)}_{0}$ to $f^{(10)}_{0}$ for each test instance and visualize this against the standard deviation across $\bar{f}_{1}$ to $\bar{f}_{10}$ in Fig. 2(a)(right).
The spread within a deep ensemble does not correlate as well with the spread across redraws of the data as the bootstrap spread does. The spread within a deep ensemble is often significantly higher than spread under redraws of the data. This illustrates an important but often ignored point—variability in ensembled predictions is not the same as variability within the ensemble. If practitioners are interested in estimating the frequentist variability of their ensemble prediction, it may be better to use bootstrap rather than looking at variability within the ensemble.
4.2 Active Learning
Perhaps the most celebrated use of epistemic uncertainty is active learning and accordingly we study the performance of our measure for such a task. Specifically, we consider a problem setup wherein an algorithm must select points to label from a pool of unlabeled points in order to train a model with high predictive accuracy.
Experimental Details: We follow the exact same experimental procedure as in Smith et al., (2023). Specifically, we use the MNIST dataset (LeCun et al.,, 1998) and train a two layer convolutional neural network with Dropout layers. We train the model for up to 50000 steps of full batch gradient descent with early stopping based on validation set performance. We provide exact hyperparameter details in Appendix B.2.
We initialize the labeled data with just 2 samples from each of the 10 classes and form a pool dataset with 4000 samples from each class from which the algorithm can acquire labels. Active learning proceeds by training the model on the currently labeled dataset, estimating epistemic uncertainty over all the instances in the pool dataset and acquiring the label for the one with the highest epistemic uncertainty. We do this for up to 300 data acquisitions and track the test accuracy.
We use three different ways of estimating epistemic uncertainty. The most commonly used approach is to leverage the interpretation of Monte Carlo Dropout as Bayesian inference (Gal and Ghahramani,, 2016) wherein Dropout is enabled at inference time and stochastic forward passes are performed to obtain samples from the Bayesian predictive distribution. We obtain 100 such samples for our experiment. Next, we use our frequentist estimate of epistemic uncertainty by running Algorithm 1 for $B=5$. Continuing our investigation into the uncertainty quantification abilities of deep ensembles, we also use the spread within the ensemble to guide data acquisition. Specifically, we use $\mathrm{I}^{\textrm{deep ensemble}}$ (Sec. 3.4) which we estimate by using the set of model weights in the deep ensemble as $\Theta_{b}$ in Algorithm 1. We use a deep ensemble of size 5. Finally, as a baseline we consider completely random acquisition of labels.
Results: Fig. 2(b) plots the mean and shades two standard deviations across 10 replications of the experiment. All three methods of epistemic uncertainty estimation perform similarly, and much better than random acquisition. The success of Dropout in generating posterior samples for mutual information estimation for active learning is well studied and justified through its Bayesian interpretation. The performance of our bootstrap based measure being comparable to Dropout shows its practical success beyond the theoretical justification from Theorem 2. The success of deep ensembles may be expected from their demonstrated capabilities on other uncertainty quantification tasks, but a theoretical justification is lacking.
4.3 Connection with Deep Ensembles
This section is aimed at investigating the success of deep ensembles at uncertainty quantification. In particular, we show in an experiment that the bootstrap MI estimate is approximately equal to that from a deep ensemble. This happens when the dominant contribution to the variation is from the random seeds rather than from statistical variability in the sample, as formalized in Section 3.4.
Experimental Setup: We use a ResNet-18 architecture and the CIFAR-10 dataset (Krizhevsky,, 2009). We defer additional hyperparameter details to Appendix B.3. We first train a deep ensemble consisting of 10 models on the entire dataset. Then, we create 10 bootstrapped datasets and train a deep ensemble of 10 models on each. Recall the notation from Section 3.4. This set of $10\times 10$ bootstrap models allows us to compute $\mathrm{I}^{\textrm{resampling}}_{b}$ and $\mathrm{I}^{\textrm{seeds}}_{b}$ by using empirical averages to approximate the expectations in Eq. 3. We similarly calculate $\mathrm{I}^{\textrm{deep ensemble}}$ as in Sec. 4.2 using the complete data deep ensemble.
Results: Fig. 3(a)(left) shows a scatter plot of $\mathrm{I}^{\textrm{seeds}}_{b}$ versus $\mathrm{I}^{\textrm{resampling}}_{b}$. The dominant contribution to the total mutual information comes from $\mathrm{I}^{\textrm{seeds}}_{b}$ with the line of best fit (red) revealing that it is 5 times $\mathrm{I}^{\textrm{resampling}}_{b}$ on average. Fig. 3(a)(right) shows that $\mathrm{I}^{\textrm{seeds}}_{b}$ and $\mathrm{I}^{\textrm{deep ensemble}}$ correlate quite well with $\rho=0.87,r_{s}=0.90$. This is our key insight—the epistemic uncertainty calculated by considering models within a deep ensemble correctly targets the dominant contribution to our (asymptotically equivalent) estimate of the Bayesian epistemic uncertainty.
Figure 3: (a) Epistemic Uncertainty Decomposition: Scatter plot of epistemic uncertainty due to training stochasticity in the bootstrap procedure ($\mathrm{I}^{\textrm{seeds}}_{b}$) versus (left) epistemic uncertainty due to data resampling in the bootstrap procedure ($\mathrm{I}^{\textrm{resampling}}_{b}$) and (right) deep ensemble estimate of epistemic uncertainty ($\mathrm{I}^{\textrm{deep ensemble}}_{b}$). See Sec. 4.3 for details. (b) Influence Function Approximation: Estimates of epistemic uncertainty using Dropout, versus using influence functions to approximate predictions from bootstrapped datasets.
4.4 Data Attribution Methods
Next, we connect our work to data-attribution methods such as influence functions (Hampel et al.,, 2011; Koh and Liang,, 2017; Giordano et al.,, 2019) and datamodels (Ilyas et al.,, 2022; Park et al.,, 2023) which aim to identify training data points most responsible for a model’s predictions. These methods allow for estimation of predictions under perturbation of training data and can therefore be used to approximate the predictions of a model trained on bootstrapped data needed for our proposed measure. Any data attribution algorithm that yields a mapping from training data to model predictions can be leveraged to estimate the variance terms in (2), and hence give epistemic uncertainty estimates. In this section, we investigate the performance of such an MI estimate based on influence functions. See Appendix B.4 for a similar experiment with datamodels.
Experimental Setup: We train a single convolutional network called the LeNet-5 (LeCun et al.,, 1998) on MNIST data. Next, we calculate the influence functions for each data point as described in Koh and Liang, (2017) for the model parameters in the last two fully connected layers. An influence function for a training data point is an estimate of how much the model parameters change under an infinitesimally small change in the weight of the training point. For each data point, we multiply the influence function by the change in the weighting under bootstrap and add all these products to the parameters to obtain an approximation for model parameters under a bootstrap reweighting. We obtain predictions from 100 such approximated models and use Algorithm 1 to obtain MI estimates. We then use Monte Carlo Dropout to generate posterior 100 predictions as described in Sec. 4.2 for the calculation of MI.
Results: Fig. 3(b) shows the influence function approximation and Dropout estimates. While there is correlation, influence function estimates are an order of magnitude smaller. We hypothesize that this is because the influence function approximation only uses a single trained model and hence fails to capture the randomness due to training stochasticity. Nonetheless, this experiment serves as a proof-of-concept for the potential of data attribution methods in uncertainty quantification.
5 Related Work
5.1 Uncertainty Decomposition
MI as a measure of information gain has demonstrated success in active learning (Houlsby et al.,, 2011; Gal et al.,, 2017), and the decomposition we consider has demonstrated success over others in OOD and misclassification detection tasks (Kotelevskii et al.,, 2025). Beyond the specific information theoretic uncertainty decomposition considered in our work, a variety of other decompositions have also been proposed (Senge et al.,, 2014; Kendall and Gal,, 2017; Hofman et al.,, 2024; Lahlou et al.,, 2023; Schweighofer et al.,, 2023; Smith et al.,, 2025) and MI does have drawbacks (Wimmer et al.,, 2023). Indeed, epistemic and aleatoric uncertainty are often not cleanly separable and take on different meanings in different contexts (Kirchhof et al.,, 2025). Our work is not intended to propose the optimality of MI as an epistemic uncertainty assessment, but rather to simply study its properties from a frequentist perspective motivated by its widespread use, strong conceptual foundation and practical success.
5.2 Bayesian and Frequentist Equivalence
Rubin, (1981) considers bootstrap from a Bayesian lens, although this interpretation is hard to reconcile with traditional Bayesian inference since it considers priors over the data rather than the parameters. Closest to our paper is the work by Newton and Raftery, (1994) which studied the use of bootstrap as a way of generating samples from a Bayesian posterior. Our work expands upon this idea by specifically leveraging this connection to connect frequentist and Bayesian notions of uncertainty. Moreover, we operationalize this idea in a modern context with DNNs and use it to study deep ensembles. Finally, Fong et al., (2023) also connects Bayesian parameter uncertainty to frequentist uncertainty arising due to lack of data. While their work conceptually shares connections to ours, it consider different practical applications.
5.3 Deep Ensembles
Deep ensembles (Lakshminarayanan et al.,, 2017) have demonstrated strong uncertainty quantification performance over other popular uncertainty quantification methods (Ovadia et al.,, 2019; Ashukha et al.,, 2020). Several reasons behind their success have been put forth (Fort et al.,, 2020), including Bayesian ones (Wilson and Izmailov,, 2020) but rigorous frequentist perspectives are lacking. The use of bootstrap in the construction of deep ensembles has also been investigated but rejected because bagged ensembles were found to have worse predictive accuracy than ensembles over random seeds that use all data points (Lee et al.,, 2015). The use of Dirichlet weights instead of multinomial has never been investigated to the authors’ knowledge, which we believe to be empirically more suitable (Sec. 2.2).
6 Discussion
Limitations: Firstly, our estimator requires retraining the model which can be prohibitive in the modern era due to computational limitations and access restrictions. On the theoretical front, Theorem 2 only proves the asymptotic validity of our estimator. Since epistemic uncertainty fundamentally targets uncertainty due to lack of data, asymptotic guarantees can appear meaningless. To alleviate this concern, we show the success of our in a low data regime—active learning. Our theorems also rely on assumptions that can be unrealistic in practical settings. For example, correct specification of the model and identifiability of parameters would never be expected to hold for DNNs. While promising, our empirical investigations are limited and more comprehensive studies are needed to draw rigorous conclusions on practical validity and success.
Future Directions: Our proof-of-concept experiment in Sec. 4.4 proposes the idea of connecting data attribution methods to uncertainty quantification. Indeed, the variance of the influence function is also the variance of the estimator (Tsiatis,, 2007). This open the doors to studying “uncertainty attribution”—identifying data points or modalities responsible for model uncertainty—by estimating their influence function. Our investigations in Sec. 4.3 that attribute uncertainty to training data variability and optimization stochasticity are an example of this idea and we hope to investigate this more deeply in a future work.
Acknowledgements
The authors thank Dr. Dan Kluger for his kind help and insights into the proofs. A.J. and S.B. were partly supported by the ARPA-H ADAPT program.
Appendix
Appendix A Proofs
A.1 Problem Setup and Notation:
We refresh the problem setup and notation below. Consider a supervised learning $K$-class classification problem. Specifically,
  1. 1.
    We assume that the training data $\mathcal{D}_{n}=\{(X_{i},Y_{i}),i=1,...,n\}$ consists of feature-label pairs, where the features $X\in\mathcal{X}$ and the labels $Y\in\{1,...,K\}$.
  2. 2.
    We wish to evaluate the epistemic uncertainty in the prediction for the label $Y_{\mathrm{test}}$ for the features $X_{\mathrm{test}}$.
  3. 3.
    We denote our predictions by $\hat{p}(X_{\mathrm{test}};\theta)$ where $\theta\in\Theta$ our model parameters. Further, denote by $\hat{p}_{k}(X_{\mathrm{test}};\theta)$ the prediction probability assigned to the $k^{\textrm{th}}$ class. The model is specified such that $0<\hat{p}_{k}(X_{\mathrm{test}};\theta)<1$ and $\sum_{k=1}^{K}\hat{p}_{k}(X_{\mathrm{test}};\theta)=1$.
  4. 4.
    In the Bayesian case, we have a prior $p(\theta)$ over the parameters and $p(\theta|\mathcal{D}_{n})$ is our posterior. Recall that the epistemic uncertainty measure in the Bayesian case is then
    $\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})$ $=\mathrm{H}(Y_{\mathrm{test}}|X_{\mathrm{test}},\mathcal{D}_{n})-\mathrm{H}(Y_{\mathrm{test}}|\theta,X_{\mathrm{test}},\mathcal{D}_{n})$
    $=\mathrm{H}(\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}(X_{\mathrm{test}};\theta)])-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\mathrm{H}(\hat{p}(X_{\mathrm{test}};\theta))].$
  5. 5.
    In the bootstrap case, we denote our estimate of $\theta$ over the bootstrapped dataset as $\hat{\theta}_{b}$. We further denote the distribution of this estimate (induced by the randomness in the choice of bootstrap weights) given the data $\mathcal{D}_{n}$ as $p(\hat{\theta}_{b}|\mathcal{D}_{n})$. Recall that our proposed epistemic uncertainty measure is then
    $\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})=\mathrm{H}(\mathbb{E}_{\hat{\theta}_{b}\sim p_{b}(\hat{\theta}_{b}|\mathcal{D}_{n})}[\hat{p}(X_{\mathrm{test}};\hat{\theta}_{b})])-\mathbb{E}_{\hat{\theta}_{b}\sim p_{b}(\hat{\theta}_{b}|\mathcal{D}_{n})}[\mathrm{H}(\hat{p}(X_{\mathrm{test}};\hat{\theta}_{b}))].$
    The bootstrap weights may be drawn from either a symmetric Dirichlet with concentration parameter 1 or a symmetric multinomial distribution.
A.2 Assumptions
  1. (A1)
    $\Theta\subset\mathbb{R}^{p}$ is open.
  2. (A2)
    There exists a $\theta_{0}\in\Theta$ such that $(X_{i},Y_{i})\stackrel{{\scriptstyle iid}}{{\sim}}P_{X}(x)\hat{p}_{y}(x;\theta_{0})=:P_{\theta_{0}}$.
  3. (A3)
    Assume
    $\psi_{\theta}(x,y):=\frac{\partial\ln(\hat{p}_{y}(x;\theta))}{\partial\theta}\Bigg|_{\theta=\theta}$
    exists everywhere. Further assume $\mathbb{E}_{(x,y)\sim P_{\theta_{0}}}[\psi_{\theta_{0}}(x,y)]=0$.
  4. (A4)
[…]
  5. (A5)
    The prior $p(\theta)$ is absolutely continuous with respect to the Lebesgue measure in a neighborhood of $\theta_{0}$, with a continuous positive density at $\theta_{0}$.
  6. (A6)
    For every $\varepsilon>0$ there exists a sequence of tests $\phi_{n}$ such that
[…]
  7. (A7)
    We further make the following assumption which implies uniform integrability to ensure that the first three moments of the posterior distribution converge
    $\sup_{n}\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[(\sqrt{n}(\hat{p}(X_{\mathrm{test}};\theta)-\hat{p}(X_{\mathrm{test}};\theta_{0}))^{3+\delta}]<\infty\quad\textrm{for some $\delta>0$}.$
  8. (A8)
    For any sequence $\{\theta_{n}\}\in\Theta$, $\mathbb{E}_{(x,y)\sim P_{\theta_{0}}}[\psi_{\theta_{n}}(x,y)]\rightarrow 0$ implies $\theta_{n}-\theta_{0}||\rightarrow 0$.
  9. (A9)
    The class $\{\psi_{\theta}:\theta\in\Theta\}$ is strong Glivenko-Cantelli.
  10. (A10)
[…]
  11. (A11)
    Denote our parameter estimate over the complete dataset as $\hat{\theta}$. Then, $\sum_{i=1}^{n}\psi_{\hat{\theta}}(X_{i},Y_{i})=o_{p}(n^{-1/2})$. Likewise, for our bootstrap estimate, $\sum_{i=1}^{n}\xi_{i}\psi_{\hat{\theta}_{b}}(X_{i},Y_{i})=o_{p}(n^{-1/2})$ where $\xi_{i},i=1,...,n$ are the bootstrap weights.
  12. (A12)
    We again make the following assumption which implies uniform integrability to ensure that the first three moments of the bootstrap distributions converge:
    $\sup_{n}\mathbb{E}_{\hat{\theta}_{b}\sim p(\hat{\theta}_{b}|\mathcal{D}_{n})}[(\sqrt{n}(\hat{p}(X_{\mathrm{test}};\hat{\theta}_{b})-\hat{p}(X_{\mathrm{test}};\theta_{0})))^{3+\delta}]<\infty\quad\textrm{for some $\delta>0$}.$
A.3 Theorem 1
Under the assumptions (A1)-(A7), we have the asymptotic expansion
$\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})=\frac{1}{2n}\sum_{k=1}^{K}\frac{\sigma_{k}^{2}}{\hat{p}_{k}(X_{\mathrm{test}};\theta_{0})}+o_{p}(n^{-1})\qquad\forall\ X_{\mathrm{test}},$
where $\sigma_{k}^{2}$ is
$\sigma_{k}^{2}=\left[\frac{\partial\hat{p}(X_{\mathrm{test}};\theta_{0})}{\partial\theta}^{T}\mathcal{I}^{-1}(\theta_{0})\frac{\partial\hat{p}(X_{\mathrm{test}};\theta_{0})}{\partial\theta}\right]_{k,k},$
for $\mathcal{I}^{-1}(\theta_{0})$ being the inverse Fisher information of the model at $\theta_{0}$.
A.3.1 Proof
Assume that we have some prior $p(\theta)$ over our parameters and denote our posterior by $p(\theta|\mathcal{D}_{n})$. Note that our epistemic uncertainty measure is then
$\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})$ $=\mathrm{H}(Y_{\mathrm{test}}|X_{\mathrm{test}},\mathcal{D}_{n})-\mathrm{H}(Y_{\mathrm{test}}|\theta,X_{\mathrm{test}},\mathcal{D}_{n})$
$=\mathrm{H}(\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}(X_{\mathrm{test}};\theta)])-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\mathrm{H}(\hat{p}(X_{\mathrm{test}};\theta))].$
Our proof proceeds in two parts, we first use Taylor’s theorem to find an expansion of the epistemic uncertainty and then we find the asymptotic limit.
Note that $\mathrm{H}$ is infinitely differentiable with first three derivatives given by
$\frac{\partial\mathrm{H}(\hat{p})}{\partial\hat{p}_{i}}$ $=-1-\log(\hat{p}_{i}),$
$\frac{\partial^{2}\mathrm{H}(\hat{p})}{\partial\hat{p}_{i}\partial\hat{p}_{j}}$ $={\rm 1}\kern-2.40005pt{\rm I}\{i=j\}\left(\frac{-1}{\hat{p}_{i}}\right),$
$\frac{\partial^{3}\mathrm{H}(\hat{p})}{\partial\hat{p}_{i}\partial\hat{p}_{j}\partial\hat{p}_{k}}$ $={\rm 1}\kern-2.40005pt{\rm I}\{i=j=k\}\left(\frac{1}{\hat{p}_{i}^{2}}\right).$
Taylor expanding $\mathrm{H}(\hat{p}(X_{\mathrm{test}};\theta))$ around $\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}(X_{\mathrm{test}};\theta)]$ yields
$\mathrm{H}(\hat{p}(X_{\mathrm{test}};\theta))=$ $\mathrm{H}(\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}(X_{\mathrm{test}};\theta)])$
$+\sum_{k=1}^{K}(\hat{p}_{k}(X_{\mathrm{test}};\theta)-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)])(-1-\log(\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)])$
$+\frac{1}{2}\sum_{k=1}^{K}(\hat{p}_{k}(X_{\mathrm{test}};\theta)-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)])^{2}\left(\frac{-1}{\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)]}\right)$
$+\frac{1}{6}\sum_{k=1}^{K}(\hat{p}_{k}(X_{\mathrm{test}};\theta)-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)])^{3}\left(\frac{1}{\hat{p}_{k}^{\prime 2}}\right),$
where $\hat{p}_{k}^{\prime}$ is some value between $\hat{p}_{k}(X_{\mathrm{test}};\theta)$ and $\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)]$ for k=1,,Kk=1,...,K.
Substituting this back into the expression for epistemic uncertainty yields
$\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})$
$=\mathrm{H}(\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}(X_{\mathrm{test}};\theta)])$
$-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\mathrm{H}(\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}(X_{\mathrm{test}};\theta)])]$
$-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}\left[\sum_{k=1}^{K}(\hat{p}_{k}(X_{\mathrm{test}};\theta)-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)])(-1-\log(\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)])\right]$
$-\frac{1}{2}\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}\left[\sum_{k=1}^{K}(\hat{p}_{k}(X_{\mathrm{test}};\theta)-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)])^{2}\left(\frac{-1}{\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)]}\right)\right]$
$-\frac{1}{6}\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}\left[\sum_{k=1}^{K}(\hat{p}_{k}(X_{\mathrm{test}};\theta)-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)])^{3}\left(\frac{1}{\hat{p}_{k}^{\prime 2}}\right)\right].$
Note that the first and second term cancel, and the third term is equal to zero. Thus,
$\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})=$ $\frac{1}{2}\sum_{k=1}^{K}\left(\frac{\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}\left[(\hat{p}_{k}(X_{\mathrm{test}};\theta)-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)])^{2}\right]}{\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)]}\right)$
[…]
Note that the first term above is the variance of the probability predictions divided by the mean prediction under the posterior distribution, i.e.
$\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})=$ $\frac{1}{2}\sum_{k=1}^{K}\left(\frac{\mathbb{V}\textrm{ar}_{\theta\sim p(\theta|\mathcal{D}_{n})}\left[\hat{p}_{k}(X_{\mathrm{test}};\theta)\right]}{\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)]}\right)$
[…]
Next, we shall evaluate the limit.
Under assumptions  (A1)(A7), we can apply the Bernstein-von Mises theorem (van der Vaart, (2000) Theorem 10.1). Note that Bernstein von-Mises implies convergence in total variation distance which directly implies convergence in distribution. Hence, we have
$\sqrt{n}\left(\theta-\theta_{0}\right)\big|\mathcal{D}_{n}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,\mathcal{I}^{-1}(\theta_{0}))\qquad\textrm{in $P_{\theta_{0}}$-probability}.$
In particular, the above is short-hand for
$\sup_{h\in BL_{1}}\left|\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}\left[h\left(\sqrt{n}(\theta-\theta_{0}\right)\right]-\mathbb{E}_{Z\sim\mathcal{N}(0,\mathcal{I}^{-1}(\theta_{0}))}[h(Z)]\right|\stackrel{{\scriptstyle P_{\theta_{0}}}}{{\rightarrow}}0,$
where $BL_{1}$ is the space of functions $h:\mathbb{R}^{p}\mapsto\mathbb{R}$ with Lipschitz norm bounded by 1. Recall that if both expectations on the left hand side were unconditional, limit of the left hand side going to 0 is the standard definition of convergence in distribution. See van der Vaart, (2000) Sec. 23.2.1 for further discussion.
We shall henceforth continue with the $(.)\stackrel{{\scriptstyle d}}{{\rightarrow}}$ (.)  in $P_{\theta_{0}}$ probability” notation for ease of exposition.
Note that (A3) allows us to apply the functional delta method (see van der Vaart, (2000) chapter 20) to evaluate the limit of the function $\theta\mapsto\hat{p}_{k}(X_{\mathrm{test}};\theta)$ under the posterior. This yields,
$\sqrt{n}\left(\hat{p}_{k}(X_{\mathrm{test}};\theta)-\hat{p}_{k}(X_{\mathrm{test}};\theta_{0})\right)\big|\mathcal{D}_{n}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,\sigma^{2}_{k})\qquad\textrm{in $P_{\theta_{0}}$-probability},$ (4)
where $\sigma^{2}_{k}=\left[\frac{\partial\hat{p}(X_{\mathrm{test}};\theta_{0})}{\partial\theta}^{T}\mathcal{I}^{-1}(\theta_{0})\frac{\partial\hat{p}(X_{\mathrm{test}};\theta_{0})}{\partial\theta}\right]_{k,k}$. Finally, (A7) implies uniform integrability allowing us to conclude that the limit of the moment is the moment of the limiting distribution:
$\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)]\stackrel{{\scriptstyle P_{\theta_{0}}}}{{\rightarrow}}\hat{p}_{k}(X_{\mathrm{test}};\theta_{0}),$
$\mathbb{V}\textrm{ar}_{\theta\sim p(\theta|\mathcal{D}_{n})}\left[\sqrt{n}\hat{p}_{k}(X_{\mathrm{test}};\theta)\right]\stackrel{{\scriptstyle P_{\theta_{0}}}}{{\rightarrow}}\sigma_{k}^{2}.$
Next, note that $\hat{p}_{k}^{\prime}\in[\hat{p}_{k}(X_{\mathrm{test}};\theta),\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)]]$. By assumption, $0<\hat{p}_{k}(X_{\mathrm{test}};\theta)<1$ and hence $1/\hat{p}_{k}^{\prime 2}=O_{p}(1)$. Furthermore, we also have from equation 4 that $\left(\hat{p}_{k}(X_{\mathrm{test}};\theta)-\hat{p}_{k}(X_{\mathrm{test}};\theta_{0})\right)\big|\mathcal{D}_{n}=O_{p}\left(\frac{1}{\sqrt{n}}\right)$. Therefore,
$n\frac{(\hat{p}_{k}(X_{\mathrm{test}};\theta)-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)])^{3}}{\hat{p_{k}^{\prime 2}}}=nO_{p}\left(\frac{1}{n^{3/2}}\right)O_{p}(1)=O_{p}\left(\frac{1}{\sqrt{n}}\right).$
Again, by the uniform integrability assumption
$\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}\left[n\frac{(\hat{p}_{k}(X_{\mathrm{test}};\theta)-\mathbb{E}_{\theta\sim p(\theta|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\theta)])^{3}}{\hat{p_{k}^{\prime 2}}}\right]\stackrel{{\scriptstyle P_{\theta_{0}}}}{{\rightarrow}}0.$
Therefore,
$n\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})\stackrel{{\scriptstyle P_{\theta_{0}}}}{{\rightarrow}}\frac{1}{2}\sum_{k=1}^{K}\frac{\sigma_{k}^{2}}{\hat{p}_{k}(X_{\mathrm{test}};\theta_{0})}.$
A.4 Theorem 2
Under the assumptions (A1)-(A12),
$\frac{\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})}{\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})}\stackrel{{\scriptstyle P_{\theta_{0}}}}{{\rightarrow}}1\qquad\forall\ X_{\mathrm{test}},$
In particular, note that $\mathrm{I}_{b}$ may be calculated using either symmetric Dirichlet weights with concentration parameter 1, or multinomial weights.
A.4.1 Proof
Recall that our proposed epistemic uncertainty measure for the bootstrap case is
$\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})=\mathrm{H}(\mathbb{E}_{\hat{\theta}_{b}\sim p_{b}(\hat{\theta}_{b}|\mathcal{D}_{n})}[\hat{p}(X_{\mathrm{test}};\hat{\theta}_{b})])-\mathbb{E}_{\hat{\theta}_{b}\sim p_{b}(\hat{\theta}_{b}|\mathcal{D}_{n})}[\mathrm{H}(\hat{p}(X_{\mathrm{test}};\hat{\theta}_{b}))].$
We can proceed exactly as in the Bayesian case to obtain
$\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})=$ $\frac{1}{2}\sum_{k=1}^{K}\left(\frac{\mathbb{V}\textrm{ar}_{\hat{\theta}_{b}\sim p_{b}(\hat{\theta}_{b}|\mathcal{D}_{n})}\left[\hat{p}_{k}(X_{\mathrm{test}};\hat{\theta}_{b})\right]}{\mathbb{E}_{\hat{\theta}_{b}\sim p_{b}(\hat{\theta}_{b}|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\hat{\theta}_{b})]}\right)$
[…]
where $\hat{p}_{k}^{\prime}$ is now some value between $\hat{p}_{k}(X_{\mathrm{test}};\hat{\theta}_{b})$ and $\mathbb{E}_{\hat{\theta}_{b}\sim p_{b}(\hat{\theta}_{b}|\mathcal{D}_{n})}[\hat{p}_{k}(X_{\mathrm{test}};\hat{\theta}_{b})]$ for k=1,,Kk=1,...,K. Next, from Theorem 10.16 of Kosorok, (2007), we have
$\sqrt{n}\left(\hat{\theta}_{b}-\hat{\theta}\big|\mathcal{D}_{n}\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,\mathcal{I}^{-1}(\theta_{0}))\qquad\textrm{in $P_{\theta_{0}}$-probability}$
and
$\sqrt{n}\left(\hat{\theta}-\theta_{0}\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,\mathcal{I}^{-1}(\theta_{0})),$
where $\hat{\theta}$ is the MLE over the complete data. Notably, both the statements above hold for either multinomial or Dirichlet weights.
We can now apply the delta method for bootstrap (van der Vaart, (2000) Theorem 23.9) to this and proceed exactly as in the Bayesian case to conclude
$n\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})\stackrel{{\scriptstyle P_{\theta_{0}}}}{{\rightarrow}}\frac{1}{2}\sum_{k=1}^{K}\frac{\sigma_{k}^{2}}{\hat{p}_{k}(X_{\mathrm{test}};\theta_{0})}.$
Combining the Bayesian and bootstrap results by the continuous mapping theorem,
$\frac{n\mathrm{I}(Y_{\mathrm{test}};\theta|X_{\mathrm{test}},\mathcal{D}_{n})}{n\mathrm{I}_{b}(X_{\mathrm{test}},\mathcal{D}_{n})}\stackrel{{\scriptstyle P_{\theta_{0}}}}{{\rightarrow}}1.$
Appendix B Additional Experiments and Experimental Details
B.1 Validation of the Plug-In Principle
Experimental Details: The original SVHN (Netzer et al.,, 2011) dataset contains 73257 training set images, 26032 test set images and 531131 extra images. We combine the original training fold and extra images fold and then use this joint set to create 11 training datasets of 50,000 images each for our experiment. We keep the test set the same.
We use a standard ResNet-18 (He et al.,, 2015) architecture but modify the first convolutional layer to have a kernel size of 3, stride of 1, padding of 1 to allow the network to work with 32 $\times$32 size images. Further, we modify the last fully connected later to only have 10 output neurons for a 10 class classification task. We use cross entropy loss and Adam (Kingma and Ba,, 2015) as an optimizer with learning rate of 0.001 and betas of 0.9 and 0.999. We use a batch size of 128 and train for 15 epochs.
The rest of the experiment proceeds as described in the main text.
B.2 Active Learning
Experimental Details: We follow the exact same setup as in Smith et al., (2023). We use the MNIST (LeCun et al.,, 1998) dataset. We first randomly select 4000 images per class to form the pool dataset. Out of these, we initialize the labeled dataset with 2 images per class and a validation dataset of 6 images per class. We keep the test set the same as in the original MNIST dataset.
We use a 2 layer convolutional network with 2 fully connected layers at the end. We use Dropout layers with probability $0.5$ between each convolutional and fully connected layer for all acquisition functions, but only for the Dropout estimate of mutual information, we keep the Dropout layers active during inference as in Gal and Ghahramani, (2016).
We train for up to 50,000 steps of full batch gradient descent with a learning rate of 0.01 and weight decay of 0.0001. We track validation log loss and after 5000 steps of non improvement, we trigger an early stopping routine which restores the model to the best validation set performance and terminates training.
The rest of the experiment proceeds as described in the main text.
B.3 Epistemic Uncertainty Decomposition
Experimental Details: We use the CIFAR-10 dataset (Krizhevsky,, 2009) with the original training and test set folds.
We use a standard ResNet-18 (He et al.,, 2015) architecture but modify the first convolutional layer to have a kernel size of 3, stride of 1, padding of 1 to allow the network to work with 32 $\times$32 size images. Further, we modify the last fully connected later to only have 10 output neurons for a 10 class classification task. We use cross entropy loss and Adam (Kingma and Ba,, 2015) as an optimizer with learning rate of 0.001 and betas of 0.9 and 0.999. We use a batch size of 128 and train for 15 epochs.
The rest of the experiment proceeds as described in the main text.
B.3.1 OOD Detection Results
We also assess the utility of our epistemic uncertainty estimates on a downstream out-of-distribution (OOD) detection task. We create 19 OOD versions of the CIFAR-10 test by applying the widely accepted corruptions proposed by Hendrycks and Dietterich, (2019) at 5 severity levels each. We then perform OOD detection on a test set created by appending the corrupted images to the original uncorrupted test set.
Table 1 lists the AUC-ROC values averaged across the 19 corruptions for each severity level for different uncertainty estimates. Specifically, $\mathrm{I}^{\textrm{deep ensemble}},\mathrm{I}^{\textrm{seeds}}_{b},\mathrm{I}^{\textrm{resampling}}_{b},\mathrm{I}_{b}$ are various mutual information estimates as described in the main text. $\mathrm{H}^{\textrm{deep ensemble}}$ is the entropy of the ensemble prediction. $\mathrm{H}_{b}$ is the entropy of the bootstrap ensemble prediction (i.e. average prediction across all bootstrap models). “Top Softmax Score” is a baseline that simply consists of using the highest class probability to order the points to form the ROC curve.
Firstly, note that $\mathrm{I}^{\textrm{resampling}}_{b}$ performs the worst perhaps surprisingly, and $\mathrm{I}^{\textrm{seeds}}_{b}$ and $\mathrm{I}_{b}$ perform very similarly as may be expected due to $\mathrm{I}^{\textrm{resampling}}_{b}$ being small and hence $\mathrm{I}_{b}\approx\mathrm{I}^{\textrm{seeds}}_{b}$. $\mathrm{I}^{\textrm{deep ensemble}}$ performs marginally better but all epistemic uncertainty measures perform worse than the simple baseline. The total uncertainty measures (based on predictive entropy) perform better with the predictive entropy of the ensemble being the best as may be expected in light of previous similar empirical findings. Furthermore, Kotelevskii et al., (2025) also observe that total entropy measures perform better for OOD detection tasks and offer discussions as to why.
Table 1: AUC-ROC for OOD detection of CIFAR-10 Corruptions
$\mathrm{I}^{\textrm{deep ensemble}}$ $\mathrm{I}^{\textrm{seeds}}_{b}$ $\mathrm{I}^{\textrm{resampling}}_{b}$ $\mathrm{I}_{b}$ $\mathrm{H}^{\textrm{deep ensemble}}$ $\mathrm{H}_{b}$ Top Softmax Score
Severity 1 0.565 0.561 0.557 0.561 0.566 0.562 0.565
Severity 2 0.613 0.606 0.599 0.606 0.615 0.608 0.613
Severity 3 0.649 0.640 0.630 0.640 0.652 0.643 0.649
Severity 4 0.682 0.674 0.661 0.673 0.687 0.678 0.683
Severity 5 0.724 0.716 0.703 0.716 0.735 0.725 0.728
B.4 Data Attribution
Experimental Details: We use the MNIST dataset. For training, we randomly select 600 images from each class to form the training dataset and keep the test dataset the same as the original.
We use a two layer convolutional neural network with 3 fully connected layers. We use Adam (Kingma and Ba,, 2015) as an optimizer with a learning rate of 0.001, betas of 0.9 and 0.999, weight decay of 0.0001, and a batch size of 32. We train for 10 epochs.
To calculate the influence function, we freeze all but the final two fully connected layers and calculate the hessian and gradients only with respect to the parameters in these layers. Specifically, as in Koh and Liang, (2017), the change in model parameters (for the final two layers) for an $\epsilon$ increase in the weighting of the $i^{th}$ training data point is $)-H_{\hat{\theta}}^{-1}\nabla_{\theta}\log(\hat{p}_{Y_{i}}(X_{i};\hat{\theta}$)) where $\hat{\theta}$ is the MLE on the original (unweighted) dataset, $)H_{\hat{\theta}}=\frac{1}{n}\sum_{i=1}^{n}\nabla^{2}_{\theta}\log(\hat{p}_{Y_{i}}(X_{i};\hat{\theta}$)) is the hessian and $\nabla_{\theta}$ denotes derivative with respect to $\theta$. We use multinomial bootstrap weights. For practical calculation, we add the identity matrix scaled by $10^{-5}$ to the Hessian to ensure invertibility in cases with non-negative definite hessian—see Koh and Liang, (2017) for discussion.
For the Dropout estimate, we add Dropout layers with probability 0.5 between the fully connected layers and train with the same hyperparameters above.
The rest of the experiment proceeds as in the main text.
B.4.1 Datamodels Experiment
Figure 4: Datamodels Approximation: Datamodel approximation to our measure versus mutual information estimates obtained using Dropout on CIFAR-10 test points. See text for details.
We now also investigate the use of datamodels (Ilyas et al.,, 2022) as a means of approximating our measure.
Experimental Details: We use the CIFAR-10 dataset (Krizhevsky,, 2009) and ResNet-9 architecture as in Ilyas et al., (2022). Each test point is associated with a datamodel which consists of a weight vector of length 50001—the number of training data points and a bias term. The datamodel is such that it predicts the model output for that test data point given the weights assigned to the training data as the dot product between the training data weights and the datamodel weights. The model output considered in Ilyas et al., (2022) is unfortunately not the entire probability vector, but simply the difference between the correct logit and the highest incorrect logit. For our experiment, we use the datamodels for CIFAR-10 provided as part of the paper—see Ilyas et al., (2022) for details. We then draw 100 sets of Dirichlet weights and obtain estimates of model outputs using the datamodel. We then pass the model output through the sigmoid function to obtain a probability and use this for calculation of the EU. Note that this is equivalent to treating the classification task as binary between the correct class vs the highest incorrect class. While this is not equivalent to the correct calculation of EU over the 10 class probability vector, it is the best we can do with the existing datamodel.
For the dropout estimate, we train a ResNet-9 with Dropout layers added with probability 0.5 with the exact same hyperparameters used to create the datamodel.
Results: Fig. 4 shows that the datamodels approximation to our measure and the Dropout epistemic uncertainty scores have a high rank correlation and moderate Pearson correlation. The datamodel estimates are also consistently lower than the Dropout estimates. This may be because of the reduction to a binary classification task in the Datamodels approximation versus the full 10 class classification in the Dropout calculation. Nonetheless, this is a useful proof of concept experiment which opens the door to using existing data attribution methods for approximate calculation of our measure.